OpenStudy (anonymous):

Factor completely 3x4 - 30x3 + 75x2. 3(x - 5)2 3x2(x - 5)2 3x2(x2 - 10x + 25) 3x2(x + 5)(x - 5)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

well i think it might be c

OpenStudy (skullpatrol):

Yes, that is the first step, but can you factor c any further since the question asks factor "completely"?

OpenStudy (anonymous):

i think so because 10 and 25 aren't simplified enough right?

OpenStudy (skullpatrol):

right, there is a relationship between them...

OpenStudy (skullpatrol):

10 = 2*(5) 25 = 5^2

OpenStudy (anonymous):

yes so its would be d

OpenStudy (anonymous):

nvmd crap

OpenStudy (skullpatrol):

$$\Huge 3x^4 - 30x^3 + 75x^2$$$$\Huge =3x^2(x^2 - 10x + 25)$$

OpenStudy (skullpatrol):

$$\Huge 3x^4 - 30x^3 + 75x^2$$$$\Huge =3x^2(x^2 - 10x + 25)$$$$\Huge a^2-2ab+b^2=(a-b)^2$$

OpenStudy (anonymous):

yep lost

OpenStudy (anonymous):

god im so failing

OpenStudy (skullpatrol):

$$\Huge x^2 - 10x + 25=x^2-2(5)x+5^2$$

OpenStudy (skullpatrol):

$$\Huge x^2 - 10x + 25= (x-5)^2$$

OpenStudy (anonymous):

then you would distribute the exponent 2 right?

OpenStudy (skullpatrol):

nope, you have the answer right there^^

OpenStudy (anonymous):

oh see im terrible at math

OpenStudy (anonymous):

so it would be d then?

OpenStudy (skullpatrol):

Be patient and put the pieces together that I gave you...

OpenStudy (anonymous):

ok give me a second

OpenStudy (anonymous):

b?

OpenStudy (skullpatrol):

yes

OpenStudy (anonymous):

omg thank you so much!!! i'm definitely giving you a medal!

OpenStudy (skullpatrol):

Thanks for trying so hard to learn :)

OpenStudy (anonymous):

thanks for helping!!